# Long-context LLMs Struggle with Long In-context Learning

### v1

@@ 1 Introduction
One line of research is based on AliBi (Press et al., 2022) and RoPE (Su et al., 2024) embedding, which allows us to train Transformers with short sequences and subsequently apply them to longer sequences during inference. Recently, different approaches (Xiong et al., 2023; Fu et al., 2024; Liu et al., 2024) help the model to extrapolate to 128K window size with continued pre-training. Later on, LongRoPE (Ding et al., 2024) was proposed to further extend to the context window to 2M tokens. Another line of research also utilizes methodologies like context window sliding and segmentation to overcome the issue of the limited context window in original Transformers (Hao et al., 2022; Ratner et al., 2023). Furthermore, architectural innovations, transitioning from traditional Transformer-based designs to recurrent models or state space models, have shown promise in facilitating long-range computations naturally Orvieto et al. (2023); Gu & Dao (2023); Peng et al. (2023a). These techniques have been incorporated into several current open-source LLMs to enhance long sequence understanding capability (Chen et al., 2023b; Tworkowski et al., 2023).
@@
Evaluations (1) and (2) only provide a minimum bar for LLMs to pass, but their results cannot reflect LLMs’ true ability to deal with realistic long-sequence tasks. Evaluation (3) provides a more realistic metric, however, these tasks are more focused on retrieving correct information from the long input. In question answering, LLMs can take shortcut to read a short snippet to predict the answer without reading the entire document as demonstrated in Figure 2 case (b). Similarly, summarization also suffers from the strong position bias, where LLMs can utilize the few leading sentences (Nallapati et al., 2017) to achieve high performance. Therefore, these metrics are insufficient to measure LLMs’ ability to comprehend and reason over the entire input sequence.
In this paper, we propose to adopt in-context learning (ICL) on extreme-label classification tasks (Anil et al., 2022; Milios et al., 2023) to evaluate long-context LLMs. Unlike the prior tasks, in-context learning requires LLMs to recognize the task by scanning over the entire input to understand the label space. This task necessitates LLMs’ ability to comprehend the entire input to make predictions. Due to the massive label space, the task demonstration could easily become a long sequence. For example, Discovery (Sileo et al., 2019) encompasses 174 classes with each example taking an average of 61 tokens. Therefore, the minimum demonstration for 1 shot/class already exceeds 10K tokens. Normally, LLMs demands more than 1 shot/class to understand the nuances of different fine-grained labels. Thus, this task becomes a natural testbed for long context understanding.
Figure 3: Results for representative models across different evaluation datasets. The performance greatly decreases as the task becomes more challenging. Some models even decays linearly w.r.t the demonstration length.
@@
We evaluate the performance of 13 long-context LLMs and find that the performance of the models uniformly dips as the task becomes more complex (e.g. requiring longer demonstration) as shown in Figure 3. Some models like Qwen and Mistral even degrade linearly w.r.t the input length. Simultaneously, most of the models can benefit from the extensive demonstration if the length is within a certain range. As the input grows longer, it either hurts or makes the performance fluctuate as shown in Figure 1. Moreover, we make further analysis on the distribution of label position to investigate the factors that take effects on the long in-context learning capability of these models. It is shown that the position distribution of instances in the prompt can dramatically influence the performance of some of the evaluated models including GPT4-turbo.
@@
- We evaluate a line of recent long-context LLMs on LongICLBench and reveal their performances with gradually changed difficulty level. Simultaneously, we find the sensitivity of some of the long-context LLMs in regard to instance position in the prompt. We hope the evaluation results can provide more insights for the improvement of the design of long-context large language models.
@@ 2 Related Work
Long In-context Learning on LLMs As pre-trained language models continue to grow in size, in-context learning (ICL) has emerged as a favored approach for addressing a wide array of tasks without the need for extensive fine-tuning  (Dong et al., 2023). A body of research has established that increasing the number of example demonstrations can enhance ICL performance  (Liu et al., 2022; Wu et al., 2023). Nonetheless, there are studies indicating that longer input prompts can actually diminish performance  (Liu et al., 2023), with the effectiveness of prior large language models (LLMs) being constrained by the maximum sequence length encountered during their training. To counter this issue, various works have introduced memory augmentation and extrapolation techniques to support ICL with an extensive set of demonstrations  (Li et al., 2023c; Wang et al., 2023).
@@
Extreme-label Classification Extreme-label Classification involves categorizing data into one of an extremely large number of labels, and finds application across a variety of real-world domains such as emotion classification from text, named entity recognition, and biological function prediction, each requiring precise differentiation among vast label spaces (Zhang et al., 2017; Sileo et al., 2019; Demszky et al., 2020; Ding et al., 2021). Existing methods to tackle Extreme-label Classification tasks range from embedding-based approaches to fine-tuned retrievals (Bhatia et al., 2015; Vulić et al., 2021), focusing on efficiently managing and leveraging the large label space. However, integrating this task with long-context large language models presents unique challenges. The sheer scale of the label space in extreme-label classification complicates the in-context learning process, where LLMs are expected to discern fine-grained differences among labels based on extensive context  (Milios et al., 2023). These challenges make the proposed LongICLBench with a range of difficulty level a good testing scenario to evaluate the capability of long-context large language models.
@@ 3 Long In-context Evaluation
TecRED Relation Extraction 41 80 [4K, 18K]
@@ 3.2 Model and Experimental Setup
We construct a prompt following the template as shown in A.2 for each of the datasets. To fairly evaluate the open-source and API-based models with a series of input lengths, we sample the same example set for all the models with labels distributed evenly to ensure an unbiased distribution for the in-context demonstration. For instance, an input of one round will include one set of examples traversing all the types, and 5 rounds will contain instances from each of the labels for 5 times. For testing, we sample 500 examples from the test set of each dataset, simultaneously ensuring an even distribution in terms of the type of labels. All the open-source models are loaded from the weights in HuggingFace [^1] https://huggingface.co, while the API-based models are called with the scripts in the official documentations  [^2] https://platform.openai.com/docs/guides/text-generation/chat-completions-api, https://cloud.google.com/vertex-ai/generative-ai/docs/multimodal/overview.
@@ 3.3 Experiment Result
The main evaluation results are demonstrated in Table 4, Table 4, Table 6 and Table 6. For the entity recognition and relationship extraction dataset, we use the F1 score as the evaluation metric, and Accuracy is utilized for the other datasets. From the presented results, generally, we can find that models of Transformer-based architecture perform consistently better than the RNN-based ones in all the evaluated datasets. However, both of them are still falling behind the powerful API-based models, especially GPT4-turbo. For a relatively simple task like BANKING77, whose context length from 1 round to 5 rounds is 2K to 14 K, most of the models can benefit from the extensive context with more demonstrations. As shown in Figure 1 and Table 4, from 2K to 4K, there is either a huge increase nearly doubling the accuracy, or a complete failure for most of the open-source models. After 3 rounds, limited performance gain can be achieved by adding more examples. When it comes to more complicated tasks like TacRED and DialogueRE in Table 4 and Table 6, which are more urgently requiring the capability of long-context comprehension, the overall performance of all the few-shot models drops compared to BANKING77. As shown in Figure 1, except GPT4-turbo can consistently benefit from more demonstrations, all of the other models reach their peak at the middle with context length around 20K.
@@ 4.1 Scattered Distribution
In our investigation, we conducte pilot experiments on TacRED, a medium-complexity dataset, with each label type demonstrated three times, resulting in a total of 123 distinct instances (calculated as $41\times 3$). Within these experiments, instances bearing the same labels are distributed randomly to form a scattered configuration. For each instance, we track its relative position within the prompt alongside its corresponding label, thereafter computing the accuracy for each label class. As illustrated in the first row of Figure 4, the visualization delineates the accuracy of each label, aligned with its position within the prompt, where diverse colors symbolize various label types. In scenarios where class instances are scattered, certain models, such as InternLM2-7B-base, demonstrate acceptable performances—approximately 60% accuracy merely on specific labels, as highlighted by a red circle in Figure 4, regardless of the instance placements. Conversely, other models, like ChatGLM3-6B-32K, exhibit robust performance across a broad spectrum of labels. Remarkably, the GPT4-turbo model consistently surpasses an 80% accuracy threshold for the majority of label types, with only a minimal count of exceptions.
@@ 4.2 Grouped Distribution
To facilitate a clear comparison between scattered and grouped distributions, we organize instances of the same class to be adjacent within the demonstration prompts. The impact of this reorganization on model performance, both pre and post-grouping, is presented in Table 7. A pronounced trend emerges, highlighting a general decline in performance across most models after grouping instances by class. Notably, models such as Mistral-7B-v0.2-base and InternLM2-7B-base exhibit significant performance drops, underscoring a pronounced sensitivity to instance grouping. In an effort to delve deeper into this phenomenon, we visualize the accuracy of grouped labels in relation to their positions within the prompt, as illustrated in Figure 4. This visualization reveals that instances of the same class, denoted by dots of the same color, are positioned in close proximity. It became evident that some models, like InternLM2-7B-base, demonstrate high sensitivity to the distribution of instances, only handling instances with labels positioned at the end of the prompt. Conversely, other open-source models such as ChatGLM3-6B-32K, with a modest 3.3% drop in accuracy, proved to be more resilient to changes in instance positioning, maintaining high performance across varied positions. Surprisingly, even the GPT4-turbo is not immune to the challenges posed by grouped distributions, experiencing a notable decline in performance by 20.3%. This observed decrease in performance is consistent across models, unaffected by the specific positions of the labels within the prompt.
@@ 5 Conclusion
In summary, our research explores the capability of large language models on long in-context learning tasks, particularly in extreme-label classification scenarios. We curate a dataset LongICLBench consisting of long in-context learning tasks with different difficulty level in respect to the context length. Through our study, we have discovered that while LLMs show promising performance on inputs up to 20K tokens, their ability to process and understand longer sequences significantly decreases. Our exploratory experiments further highlight the impact of the distribution of examples within prompts on model performance. We hope LongICLBench and our findings contribute to the ongoing efforts to enhance LLMs’ understanding of long contexts.
@@ A.2 Prompting Template
TecRED Given a sentence and a pair of subject and object entities within the sentence, please predict the relation between the given entities. The examples are as follows:
@@
The prompting template for each of the dataset is presented at Table 10

### v2

@@ 1 Introduction
One line of research is based on AliBi (Press et al., 2022) and RoPE (Su et al., 2024) embedding, which allows us to train Transformers with short sequences and subsequently apply them to longer sequences during inference. Recently, different approaches (Xiong et al., 2023; Fu et al., 2024; Liu et al., 2024) help the model to extrapolate to 128K window size with continued pre-training. Later on, LongRoPE (Ding et al., 2024) was proposed to further extend the context window to 2M tokens. Another line of research also utilizes methodologies like context window sliding and segmentation to overcome the issue of the limited context window in original Transformers (Hao et al., 2022; Ratner et al., 2023). Furthermore, architectural innovations, transitioning from traditional Transformer-based designs to recurrent models or state space models, have shown promise in facilitating long-range computations naturally Orvieto et al. (2023); Gu & Dao (2023); Peng et al. (2023a). These techniques have been incorporated into several current open-source LLMs to enhance long sequence understanding capability (Chen et al., 2023b; Tworkowski et al., 2023).
@@
Evaluations (1) and (2) only provide a minimum bar for LLMs to pass, but their results cannot reflect LLMs’ true ability to deal with realistic long-sequence tasks. Evaluation (3) provides a more realistic metric, however, these tasks are more focused on retrieving correct information from the long input. In question answering, LLMs can take a shortcut to read a short snippet to predict the answer without reading the entire document as demonstrated in Figure 2 case (b). Similarly, summarization also suffers from the strong position bias, where LLMs can utilize the few leading sentences (Nallapati et al., 2017) to achieve high performance. Therefore, these metrics are insufficient to measure LLMs’ ability to comprehend and reason over the entire input sequence.
In this paper, we propose to adopt in-context learning (ICL) on extreme-label classification tasks (Anil et al., 2022; Milios et al., 2023) to evaluate long-context LLMs. Unlike the prior tasks, in-context learning requires LLMs to recognize the task by scanning over the entire input to understand the label space. This task necessitates LLMs’ ability to comprehend the entire input to make predictions. Due to the massive label space, the task demonstration could easily become a long sequence. For example, Discovery (Sileo et al., 2019) encompasses 174 classes with each example taking an average of 61 tokens. Therefore, the minimum demonstration for 1 shot/class already exceeds 10K tokens. Normally, LLMs demand more than 1 shot/class to understand the nuances of different fine-grained labels. Thus, this task becomes a natural testbed for long-context understanding.
Figure 3: Results for representative models across different evaluation datasets. The performance greatly decreases as the task becomes more challenging. Some models even decay linearly w.r.t the demonstration length.
@@
We evaluate the performance of 13 long-context LLMs and find that the performance of the models uniformly dips as the task becomes more complex (e.g. requiring longer demonstration) as shown in Figure 3. Some models like Qwen and Mistral even degrade linearly w.r.t the input length. Simultaneously, most of the models can benefit from the extensive demonstration if the length is within a certain range. As the input grows longer, it either hurts or makes the performance fluctuate as shown in Figure 1. Moreover, we make further analysis on the distribution of label position to investigate the factors that affect the long in-context learning capability of these models. It is shown that the position distribution of instances in the prompt can dramatically influence the performance of some of the evaluated models including GPT4-turbo.
@@
- We evaluate a line of recent long-context LLMs on LongICLBench and reveal their performances with gradually changed difficulty levels. Simultaneously, we find the sensitivity of some of the long-context LLMs regarding instance position in the prompt. We hope the evaluation results can provide more insights for the improvement of the design of long-context large language models.
@@ 2 Related Work
Long In-context Learning on LLMs As pre-trained language models continue to grow in size, in-context learning (ICL) has emerged as a favored approach for addressing a wide array of tasks without the need for extensive fine-tuning  (Dong et al., 2023). A body of research has established that increasing the number of example demonstrations can enhance ICL performance  (Liu et al., 2022; Wu et al., 2023). Nonetheless, there are studies indicating that longer input prompts can actually diminish performance  (Liu et al., 2023), with the effectiveness of prior large language models (LLMs) being constrained by the maximum sequence length encountered during their training. It is also claimed in previous works that LLM+ICL falls short on specification-heavy tasks due to inadequate long-text understanding ability  (Peng et al., 2023c). To counter this issue, various works have introduced memory augmentation and extrapolation techniques to support ICL with an extensive set of demonstrations  (Li et al., 2023c; Wang et al., 2023).
@@
Extreme-label Classification Extreme-label Classification involves categorizing data into one of an extremely large number of labels, and finds application across a variety of real-world domains such as emotion classification from text, named entity recognition, and biological function prediction, each requiring precise differentiation among vast label spaces (Zhang et al., 2017; Sileo et al., 2019; Demszky et al., 2020; Ding et al., 2021). Existing methods to tackle Extreme-label Classification tasks range from embedding-based approaches to fine-tuned retrievals (Bhatia et al., 2015; Vulić et al., 2021), focusing on efficiently managing and leveraging the large label space. However, integrating this task with long-context large language models presents unique challenges. The sheer scale of the label space in extreme-label classification complicates the in-context learning process, where LLMs are expected to discern fine-grained differences among labels based on extensive context  (Milios et al., 2023). These challenges make the proposed LongICLBench with a range of difficulty levels a good testing scenario to evaluate the capability of long-context large language models.
@@ 3 Long In-context Evaluation
TacRED Relation Extraction 41 80 [4K, 18K]
@@ 3.2 Model and Experimental Setup
We construct a prompt following the template as shown in A.2 for each of the datasets. To fairly evaluate the open-source and API-based models with a series of input lengths, we sample the same example set for all the models with labels distributed evenly to ensure an unbiased distribution for the in-context demonstration. For instance, an input of one round will include one set of examples traversing all the types, and 5 rounds will contain instances from each of the labels 5 times. For testing, we sample 500 examples from the test set of each dataset, simultaneously ensuring an even distribution in terms of the type of labels. All the open-source models are loaded from the weights in HuggingFace [^1] https://huggingface.co, while the API-based models are called with the scripts in the official documentations  [^2] https://platform.openai.com/docs/guides/text-generation/chat-completions-api, https://cloud.google.com/vertex-ai/generative-ai/docs/multimodal/overview.
@@ 3.3 Experiment Result
The main evaluation results are demonstrated in Table 4, Table 4, Table 6 and Table 6. For the entity recognition and relationship extraction dataset, we use the F1 score as the evaluation metric, and Accuracy is utilized for the other datasets. From the presented results, generally, we can find that models of Transformer-based architecture perform consistently better than the RNN-based ones in all the evaluated datasets. However, both of them are still falling behind the powerful API-based models, especially GPT4-turbo. For a relatively simple task like BANKING77, whose context length from 1 round to 5 rounds is 2K to 14 K, most of the models can benefit from the extensive context with more demonstrations. As shown in Figure 1 and Table 4, from 2K to 4K, there is either a huge increase nearly doubling the accuracy, or a complete failure for most of the open-source models. After 3 rounds, limited performance gain can be achieved by adding more examples. When it comes to more complicated tasks like TacRED and DialogueRE in Table 4 and Table 6, which are more urgently requiring the capability of long-context comprehension, the overall performance of all the few-shot models drops compared to BANKING77. As shown in the middle plot of Figure 1, only GPT4-turbo can consistently benefit from more demonstrations, all of the other models reach their peak at the middle with context length around 20K.
@@ 4.1 Scattered Distribution
In our investigation, we conducted pilot experiments on TacRED, a medium-complexity dataset, with each label type demonstrated three times, resulting in a total of 123 distinct instances (calculated as $41\times 3$). Within these experiments, instances bearing the same labels are distributed randomly to form a scattered configuration. For each instance, we track its relative position within the prompt alongside its corresponding label, thereafter computing the accuracy for each label class. As illustrated in the first row of Figure 4, the visualization delineates the accuracy of each label, aligned with its position within the prompt, where diverse colors symbolize various label types. In scenarios where class instances are scattered, certain models, such as InternLM2-7B-base, demonstrate acceptable performances—approximately 60% accuracy merely on specific labels, as highlighted by a red circle in Figure 4, regardless of the instance placements. Conversely, other models, like ChatGLM3-6B-32K, exhibit robust performance across a broad spectrum of labels. Remarkably, the GPT4-turbo model consistently surpasses an 80% accuracy threshold for the majority of label types, with only a minimal count of exceptions.
@@ 4.2 Grouped Distribution
To facilitate a clear comparison between scattered and grouped distributions, we organize instances of the same class to be adjacent within the demonstration prompts. The impact of this reorganization on model performance, both pre and post-grouping, is presented in Table 7. A pronounced trend emerges, highlighting a general decline in performance across most models after grouping instances by class. Notably, models such as Mistral-7B-v0.2-base and InternLM2-7B-base exhibit significant performance drops, underscoring a pronounced sensitivity to instance grouping. In an effort to delve deeper into this phenomenon, we visualize the accuracy of grouped labels in relation to their positions within the prompt, as illustrated in Figure 4. This visualization reveals that instances of the same class, denoted by dots of the same color, are positioned nearby. It became evident that some models, like InternLM2-7B-base, demonstrate high sensitivity to the distribution of instances, only handling instances with labels positioned at the end of the prompt. Conversely, other open-source models such as ChatGLM3-6B-32K, with a modest 3.3% drop in accuracy, proved to be more resilient to changes in instance positioning, maintaining high performance across varied positions. Surprisingly, even the GPT4-turbo is not immune to the challenges posed by grouped distributions, experiencing a notable decline in performance by 20.3%. This observed decrease in performance is consistent across models, unaffected by the specific positions of the labels within the prompt.
@@ 5 Conclusion
In summary, our research explores the capability of large language models on long in-context learning tasks, particularly in extreme-label classification scenarios. We curate a dataset LongICLBench consisting of long in-context learning tasks with different difficulty levels with respect to the context length. Through our study, we have discovered that while LLMs show promising performance on inputs up to 20K tokens, their ability to process and understand longer sequences significantly decreases. Our exploratory experiments further highlight the impact of the distribution of examples within prompts on model performance. We hope LongICLBench and our findings contribute to the ongoing efforts to enhance LLMs’ understanding of long contexts.
@@ A.2 Prompting Template
TacRED Given a sentence and a pair of subject and object entities within the sentence, please predict the relation between the given entities. The examples are as follows:
@@
The prompting template for each of the datasets is presented at Table 10